AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 6 of 33  ELOG logo
ID Date Author Subject
  563   Thu May 30 09:30:33 2024 CB, PW, YL, TDRun 10

10:30

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

 

10.48 FEE64 temperatures OK - attachment 1

WR timestamps OK - attachment 2

ADC data item stats - attachment 3

per FEE64 Rate spectra - attachment 4

 why do we observe HEC data in aida15 asic #1?

per FEE64 1.8.L spectra - attachments 5-6

 ADC to 20MeV FSR - not clear when these spwectra were last zero;d

 currently 750keV/u D + no target ?

10.53 All histograms & stats zero'd

Merger, Taper Server etc - attachments 7-8

  bind error 98 for NewMerger server - not clear when these messahes were generated - added some returns to monitor whether these, or other, messages are repeated

DSSSD bias & leakage current OK - attachment 9

11:30

We noticed some odd peak structure in the rate energy histograms for the right hand side detectors. We took away the beam but this structure persisted. 

This was not observed in the total energy histograms, the reason for this is the applied multiplicity cut to the total histograms. Removing this cut the peaks are clearly visible (attachment 10). Putting the cut back in we see the peaks removed (attachment 11). These events correspond to some noisy strips on the detectors on the right hand side (see attachment 12). Notice the much larger rate on the RHS. Current ADC data items statistics show FEE modules for the RHS also much larger (attachment 13). We did not observe this with no beam (https://elog.ph.ed.ac.uk/CARME/554) but did with beam last night (https://elog.ph.ed.ac.uk/CARME/558) so there nust be some effect of having beam in the ring increasing the number of events on the RHS. Note the detectors are fully out and the events are localised in these strips. 

Note we have moved the detectorsand removed beam from the ring in this time period to investigate these peaks.

12:50

Run 10 continues. Temps OK. System checks OK. Merger OK, Tape server OK.

Leakage current OK.

Notice the 'beam peak' at 1.5 MeV is larger in the left hand side (attachment 14)

14:26

We have started the new_optimeren.py on pi@gas-stepper to try and align the gas target. Valves to the interaction chamber are closed.

Setting the nozzle X=-13000, Y=1000 we observe an increase in E4 from 4E-10 mbar to 2E-9 mbar. Moving either side we observe a drop in E4. We take this as the gas target being aligned.

Anton will now open the valves to the gas target.

Run 10 stopped.         

Attachment 1: Screenshot_from_2024-05-30_10-48-43.png
Screenshot_from_2024-05-30_10-48-43.png
Attachment 2: Screenshot_from_2024-05-30_10-49-43.png
Screenshot_from_2024-05-30_10-49-43.png
Attachment 3: Screenshot_from_2024-05-30_10-50-34.png
Screenshot_from_2024-05-30_10-50-34.png
Attachment 4: Screenshot_from_2024-05-30_10-51-40.png
Screenshot_from_2024-05-30_10-51-40.png
Attachment 5: Screenshot_from_2024-05-30_10-52-51.png
Screenshot_from_2024-05-30_10-52-51.png
Attachment 6: Screenshot_from_2024-05-30_10-53-53.png
Screenshot_from_2024-05-30_10-53-53.png
Attachment 7: Screenshot_from_2024-05-30_10-56-23.png
Screenshot_from_2024-05-30_10-56-23.png
Attachment 8: Screenshot_from_2024-05-30_10-56-42.png
Screenshot_from_2024-05-30_10-56-42.png
Attachment 9: Screenshot_from_2024-05-30_10-57-56.png
Screenshot_from_2024-05-30_10-57-56.png
Attachment 10: Screenshot_from_2024-05-30_11-35-08.png
Screenshot_from_2024-05-30_11-35-08.png
Attachment 11: Screenshot_from_2024-05-30_11-36-07.png
Screenshot_from_2024-05-30_11-36-07.png
Attachment 12: Screenshot_from_2024-05-30_12-58-20.png
Screenshot_from_2024-05-30_12-58-20.png
Attachment 13: Screenshot_from_2024-05-30_13-10-18.png
Screenshot_from_2024-05-30_13-10-18.png
  562   Thu May 30 08:33:21 2024 CB, PW, YLRun 8 & 9- detector motion tests

09:34 - Moved to Run 9 in preparation of moving detectors. No target yet. Beam in the ring at 750 keV/A.

Temperatures OK (attach)

System wide checks OK except 16 fails clock and ADC calib

Rates OK (attach)

Writing to disk OK (attach)

Resetting the sorting script. Moving detectors.

Moved detectors in. See attached. Stripe / shadow quite clear. Some oddities on the top-right DSSD. May be due to ExEy cuts and multiplicity cuts.

Moved on tun run 9 to acquire some stats without motion.

 

10:13 Temperatures still OK. Rates OK. Leakage currents stable.

We were cutting the beam.

Increased repetition to x5.

Motors fully out - run 10 started.

Attachment 1: Screenshot_from_2024-05-30_09-33-55.png
Screenshot_from_2024-05-30_09-33-55.png
Attachment 2: Screenshot_from_2024-05-30_09-35-41.png
Screenshot_from_2024-05-30_09-35-41.png
Attachment 3: Screenshot_from_2024-05-30_09-36-23.png
Screenshot_from_2024-05-30_09-36-23.png
Attachment 4: Screenshot_from_2024-05-30_09-50-27.png
Screenshot_from_2024-05-30_09-50-27.png
Attachment 5: Screenshot_from_2024-05-30_09-52-29.png
Screenshot_from_2024-05-30_09-52-29.png
Attachment 6: Screenshot_from_2024-05-30_09-57-27.png
Screenshot_from_2024-05-30_09-57-27.png
Attachment 7: Screenshot_from_2024-05-30_10-00-25.png
Screenshot_from_2024-05-30_10-00-25.png
  561   Thu May 30 07:25:53 2024 Yuri, Phil, NikosGas target

The cave is closed and beam is back

The d intensity is about half of the intensity before the break

The gas target is prepared but the valvas cannot be opened. Anton is coming in at around 1 PM to open the gate valves.

  560   Thu May 30 06:00:51 2024 Yuri Nikos ...Open the cave

Nikos needs to enter the cave to work on the target cooling.

 

  559   Wed May 29 17:42:06 2024 TDOffline analysis APR24/R2 (pulser walkthrough)
ADC offsets using pulser walkthrough data from data file APR24/R2

ch = channel + ( module * 64 ) + ( range * 2048 )

adc_data( ch ) = INT( RSHIFT( ABS( adc_data( ch ) - 32768 ), 3 ) - offset( ch ) + 0.5 )




per FEE64 ADC spectra - attachments 1-16

Note common x & y-scales - peak height proxy for peak width

aida01 1.8.L (s7) pulser peak width 9 ch FWHM (~50keV FWHM)




ADC offset analysis

/home/td/Calibration/offsets /home/td/CARME/R2 50 2048 20 5

First pass sort variables NAMELIST I/O format ( ADC offsets, FEE64 configuration, LEC & HEC equal energy cuts ) - attachment 17
Attachment 1: Screenshot_from_2024-05-28_20-54-05.png
Screenshot_from_2024-05-28_20-54-05.png
Attachment 2: Screenshot_from_2024-05-28_20-54-58.png
Screenshot_from_2024-05-28_20-54-58.png
Attachment 3: Screenshot_from_2024-05-28_20-55-45.png
Screenshot_from_2024-05-28_20-55-45.png
Attachment 4: Screenshot_from_2024-05-28_21-14-54.png
Screenshot_from_2024-05-28_21-14-54.png
Attachment 5: Screenshot_from_2024-05-28_21-15-49.png
Screenshot_from_2024-05-28_21-15-49.png
Attachment 6: Screenshot_from_2024-05-28_21-16-20.png
Screenshot_from_2024-05-28_21-16-20.png
Attachment 7: Screenshot_from_2024-05-28_21-16-46.png
Screenshot_from_2024-05-28_21-16-46.png
Attachment 8: Screenshot_from_2024-05-28_21-18-21.png
Screenshot_from_2024-05-28_21-18-21.png
Attachment 9: Screenshot_from_2024-05-28_21-27-03.png
Screenshot_from_2024-05-28_21-27-03.png
Attachment 10: Screenshot_from_2024-05-28_21-27-42.png
Screenshot_from_2024-05-28_21-27-42.png
Attachment 11: Screenshot_from_2024-05-28_21-28-12.png
Screenshot_from_2024-05-28_21-28-12.png
Attachment 12: Screenshot_from_2024-05-28_21-28-46.png
Screenshot_from_2024-05-28_21-28-46.png
Attachment 13: Screenshot_from_2024-05-28_21-29-16.png
Screenshot_from_2024-05-28_21-29-16.png
Attachment 14: Screenshot_from_2024-05-28_21-29-40.png
Screenshot_from_2024-05-28_21-29-40.png
Attachment 15: Screenshot_from_2024-05-28_21-30-14.png
Screenshot_from_2024-05-28_21-30-14.png
Attachment 16: Screenshot_from_2024-05-28_21-30-48.png
Screenshot_from_2024-05-28_21-30-48.png
Attachment 17: variables.dat
  $variables
  offset(   0) =    31.70
  offset(   1) =     1.56
  offset(   2) =   -20.38
  offset(   3) =   -19.76
  offset(   4) =    13.07
  offset(   5) =     4.85
  offset(   6) =   -21.16
  offset(   7) =    22.27
  offset(   8) =    -5.26
  offset(   9) =    -4.43
  offset(  10) =     1.25
  offset(  11) =    -0.48
  offset(  12) =     9.68
  offset(  13) =     2.46
  offset(  14) =    -5.27
  offset(  15) =     9.41
  offset(  16) =     0.90
  offset(  17) =    -2.60
  offset(  18) =    22.25
  offset(  19) =   -19.17
  offset(  20) =     1.48
  offset(  21) =    -9.92
  offset(  22) =    11.92
  offset(  23) =    12.82
  offset(  24) =    12.26
  offset(  25) =    -7.54
  offset(  26) =    12.66
  offset(  27) =   -23.36
  offset(  28) =     9.50
  offset(  29) =    17.75
  offset(  30) =   -23.58
  offset(  31) =    21.56
  offset(  32) =     7.89
  offset(  33) =   -11.31
  offset(  34) =   -25.22
  offset(  35) =    -5.51
  offset(  36) =   -13.56
  offset(  37) =    -3.33
  offset(  38) =   -15.50
  offset(  39) =     0.09
  offset(  40) =   -12.56
  offset(  41) =   -26.14
  offset(  42) =     3.57
  offset(  43) =   -14.70
  offset(  44) =    12.66
  offset(  45) =   -36.78
  offset(  46) =    24.58
  offset(  47) =    -8.98
  offset(  48) =   -18.05
  offset(  49) =    -7.85
  offset(  50) =   -25.85
  offset(  51) =   -15.22
  offset(  52) =   -10.96
  offset(  53) =   -17.87
  offset(  54) =   -29.36
  offset(  55) =    -9.94
  offset(  56) =    10.20
  offset(  57) =   -23.00
  offset(  58) =   -22.69
  offset(  59) =   -21.56
  offset(  60) =    27.94
  offset(  61) =     3.41
  offset(  62) =     6.03
  offset(  63) =   -10.52
  offset(  64) =     7.65
  offset(  65) =     5.94
  offset(  66) =     0.52
  offset(  67) =     1.01
  offset(  68) =    -7.82
  offset(  69) =   -22.59
  offset(  70) =     9.05
  offset(  71) =    23.35
  offset(  72) =    23.79
  offset(  73) =     3.42
  offset(  74) =     2.83
  offset(  75) =     2.65
  offset(  76) =     3.05
  offset(  77) =     6.05
  offset(  78) =   -12.53
  offset(  79) =   -13.37
  offset(  80) =     5.99
  offset(  81) =    -6.94
  offset(  82) =     0.39
  offset(  83) =   -12.79
  offset(  84) =    -1.48
  offset(  85) =    20.06
  offset(  86) =     1.26
  offset(  87) =   -12.90
  offset(  88) =    18.92
  offset(  89) =    -3.11
  offset(  90) =    -8.23
  offset(  91) =    -6.89
  offset(  92) =    -7.46
  offset(  93) =     0.15
  offset(  94) =   -17.01
  offset(  95) =    -8.25
  offset(  96) =   -31.52
  offset(  97) =    16.13
  offset(  98) =    15.45
  offset(  99) =    24.67
  offset( 100) =   -20.48
  offset( 101) =     2.01
  offset( 102) =    -2.56
  offset( 103) =    42.23
  offset( 104) =     4.75
  offset( 105) =    15.76
  offset( 106) =    11.44
  offset( 107) =   -34.81
  offset( 108) =   -28.62
  offset( 109) =   -21.31
  offset( 110) =   -24.33
  offset( 111) =    -5.05
  offset( 112) =    15.62
  offset( 113) =     5.20
  offset( 114) =    42.22
  offset( 115) =    16.79
  offset( 116) =    -0.78
  offset( 117) =   -12.94
  offset( 118) =    24.25
  offset( 119) =    16.62
  offset( 120) =    -6.90
  offset( 121) =    28.25
  offset( 122) =    23.14
  offset( 123) =    11.89
  offset( 124) =    -6.48
  offset( 125) =    10.15
  offset( 126) =    55.36
  offset( 127) =    21.89
  offset( 128) =   -14.38
  offset( 129) =    -4.43
  offset( 130) =   -17.38
  offset( 131) =    35.78
  offset( 132) =   -42.46
  offset( 133) =    -8.70
  offset( 134) =    15.69
  offset( 135) =   -26.25
  offset( 136) =   -21.07
  offset( 137) =   -35.26
  offset( 138) =   -10.76
  offset( 139) =   -31.95
  offset( 140) =     6.11
  offset( 141) =   -19.00
  offset( 142) =   -13.84
  offset( 143) =   -17.20
  offset( 144) =     7.83
  offset( 145) =    -8.38
  offset( 146) =   -15.68
  offset( 147) =    -5.98
  offset( 148) =     9.85
  offset( 149) =   -21.39
  offset( 150) =   -42.83
  offset( 151) =   -55.99
  offset( 152) =    -8.53
  offset( 153) =     0.97
  offset( 154) =   -29.26
  offset( 155) =   -16.40
  offset( 156) =    -3.32
  offset( 157) =   -42.87
  offset( 158) =   -42.72
  offset( 159) =   -26.85
  offset( 160) =   -36.87
  offset( 161) =   -33.16
  offset( 162) =   -10.69
  offset( 163) =   -23.53
  offset( 164) =    10.43
  offset( 165) =   -10.48
  offset( 166) =   -31.70
  offset( 167) =    21.74
  offset( 168) =     2.27
  offset( 169) =    -7.56
  offset( 170) =   -19.47
  offset( 171) =    25.12
  offset( 172) =   -17.19
  offset( 173) =    -7.30
  offset( 174) =   -44.54
  offset( 175) =   -17.26
  offset( 176) =   -29.31
  offset( 177) =   -12.46
  offset( 178) =   -33.61
  offset( 179) = -9999.99
  offset( 180) =   -31.25
  offset( 181) =    -5.14
  offset( 182) =     5.87
  offset( 183) =   -25.05
  offset( 184) =   -34.61
  offset( 185) =   -19.87
  offset( 186) =    -3.71
  offset( 187) =   -30.59
  offset( 188) =     1.07
  offset( 189) =   -48.50
  offset( 190) =   -34.50
  offset( 191) =   -25.18
  offset( 192) =   -37.17
  offset( 193) =   -40.22
  offset( 194) =   -38.76
  offset( 195) =   -52.65
  offset( 196) =   -37.83
  offset( 197) =    -5.82
  offset( 198) =    -7.01
  offset( 199) =    -4.33
  offset( 200) =    -4.06
  offset( 201) =   -36.15
  offset( 202) =   -14.17
  offset( 203) =   -32.62
  offset( 204) =   -39.21
  offset( 205) =    -7.39
  offset( 206) =   -10.53
  offset( 207) =   -36.56
  offset( 208) =   -34.01
  offset( 209) =   -61.07
  offset( 210) =   -15.10
  offset( 211) =   -35.46
  offset( 212) =    -7.47
  offset( 213) =   -35.09
  offset( 214) =   -78.70
  offset( 215) =   -29.20
  offset( 216) =   -22.09
  offset( 217) =   -22.15
  offset( 218) =   -16.24
  offset( 219) =   -31.62
  offset( 220) =   -32.45
  offset( 221) =   -45.59
  offset( 222) =   -45.27
  offset( 223) =    10.54
  offset( 224) =    10.90
  offset( 225) =   -16.94
  offset( 226) =   -15.20
  offset( 227) =    -2.58
  offset( 228) =   -16.18
  offset( 229) =   -13.20
  offset( 230) =   -54.53
  offset( 231) =   -17.51
  offset( 232) =   -29.77
  offset( 233) =   -30.14
  offset( 234) =   -42.68
  offset( 235) =   -36.03
  offset( 236) =   -17.08
  offset( 237) =   -14.88
  offset( 238) =   -39.70
  offset( 239) =   -64.37
  offset( 240) =   -34.07
  offset( 241) =    16.67
  offset( 242) =     4.33
  offset( 243) =   -11.05
  offset( 244) =    11.48
  offset( 245) =   -30.51
  offset( 246) =    12.23
  offset( 247) =   -12.35
  offset( 248) =    14.03
  offset( 249) =   -27.00
  offset( 250) =     1.20
  offset( 251) =   -24.70
  offset( 252) =   -15.46
  offset( 253) =   -18.55
  offset( 254) =     6.72
  offset( 255) =    -7.39
  offset( 256) = -9999.99
  offset( 257) = -9999.99
  offset( 258) =    29.33
  offset( 259) =    21.68
  offset( 260) = -9999.99
  offset( 261) =    33.37
  offset( 262) =     1.22
  offset( 263) =    30.59
  offset( 264) =     2.99
  offset( 265) =    10.75
  offset( 266) =   -27.39
  offset( 267) = -9999.99
  offset( 268) =    -4.27
  offset( 269) =   -12.65
  offset( 270) =     7.85
  offset( 271) =    -5.75
  offset( 272) =   -12.38
  offset( 273) =    -0.29
  offset( 274) = -9999.99
  offset( 275) =   -39.58
  offset( 276) =   -26.97
  offset( 277) =   -19.65
  offset( 278) =   -28.28
  offset( 279) =   -23.94
  offset( 280) =     5.51
  offset( 281) = -9999.99
  offset( 282) =    -1.87
  offset( 283) =   -34.61
  offset( 284) =     6.36
  offset( 285) =   -21.14
  offset( 286) =   -24.61
  offset( 287) =   -25.32
  offset( 288) = -9999.99
  offset( 289) =    -4.38
  offset( 290) =   -13.00
  offset( 291) =    -1.66
  offset( 292) =    34.08
  offset( 293) =    -2.70
  offset( 294) =     2.43
  offset( 295) = -9999.99
  offset( 296) =     1.82
  offset( 297) =     7.79
  offset( 298) =   -27.47
... 747 more lines ...
  558   Wed May 29 16:46:21 2024 NP, AB, PMRun7 - Measurements only with ion beam.

18:19 Start or run 7: In beam has been optimzed up to now by Zoran and Claude. ~7E7 particles at injection. From now on, only the injection beamline is optimized by Operators.

Detectors IN (left arm pot =18.86 mm, right arm pot = 19.80 mm)

No Target

Beam is in the ring for ~6 seconds and total cycle time between injections is ~8 seconds

Injection beamline is still in process of optimization by operators

21:30

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

23:30

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

01:30

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

Beam and Electron Coller failure (~00:35); Beam back (~01:08); Cooler back (~01:20)

03:00

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

Beam and Electron Coller failure (~01:35); Beam back (~01:55); Cooler back (~01:55)

04:40

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

06:00

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

07:00

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

 

 

Attachment 1: Screenshot_from_2024-05-29_21-49-57.png
Screenshot_from_2024-05-29_21-49-57.png
Attachment 2: Screenshot_from_2024-05-29_21-50-14.png
Screenshot_from_2024-05-29_21-50-14.png
  557   Wed May 29 09:49:35 2024 JM, CBThreshold changes / D+ beam

09:00

R4 alpha background continues, currently R4_12

Temps OK, system checks unchanged from previous status (OK, aida16 clock and ADC issue)

10:30

R4 stopped and moved to the next run R5.

Lithium 3+ beam is not feasible at sufficient densities so we are switching to D+ at a beam energy of 0.75 MeV/u. A measurement of d+p at this energy would produce 3He at energies of ~900 keV. 

The current threshold (slow comparator) has been 1 MeV (0x64). We lowered the thresholds in all FEE modules to 200 keV (0x14). ADC rates look fine in all FEEs (~tens of Hz) except in aida 15 and aida 16 which was ~several hundred kHz. Data rate to disk was 1.2 Mb/s

We increased the threshold in aida 15 and 16 only to 500 keV (0x32) and rates in aida 15 and 16 have dropped to <20 kHz. Data rate to disk now 220 kB/s.

11:00

R5 stopped. R6 started - background with reduced thresholds. 

Attachment 1 - ADC data items statistics

Attachment 2 - Histogram browser statistics

Attachment 3 - Tape server

Attachment 4 - New Merger

13:12

R6 continues R6_9

Edits to the CARME sort code have been made. EyEx energy cut for online monitoring reduced to 100 keV. Multiplicity cut applies to lowEnergyEy, lowEnergyEx and lowEnergyXY histograms. Multiplicity cut considers only single strip events.

15:00

D+ beam is now in the ring and is being optimised by the beam operators. The CARME detectors are currently fullt out, but we can see some effects of the beam on the online monitoring code.

Attachment 1: Screenshot_from_2024-05-29_11-00-20.png
Screenshot_from_2024-05-29_11-00-20.png
Attachment 2: Screenshot_from_2024-05-29_11-01-30.png
Screenshot_from_2024-05-29_11-01-30.png
Attachment 3: Screenshot_from_2024-05-29_11-01-47.png
Screenshot_from_2024-05-29_11-01-47.png
Attachment 4: Screenshot_from_2024-05-29_11-02-03.png
Screenshot_from_2024-05-29_11-02-03.png
  556   Wed May 29 08:10:32 2024 JMRun4 - alpha background

09:00

R4 alpha background continues, currently R4_12

Temps OK, system checks unchanged from previous status (OK, aida16 clock and ADC issue)

  555   Tue May 28 15:25:21 2024 TD, JM, CBPulser tests

R3 started

Pulser ON. Approx 2 kHZ.

1.8*w waveforms attachments 1+2

ADC data items statistics - attchment 3

1.8*L spectra - attachments 4+5

Pulser widths are larger than during testing prior to the Feb beamtime. Widths in Channels;

aida01 = 76

aida02 = 82

aida03 = will upate all later tonight..

R3 ended. Pulser OFF

17:15

R4 Started. Alpha background run.

 

Attachment 1: Screenshot_from_2024-05-28_16-14-56.png
Screenshot_from_2024-05-28_16-14-56.png
Attachment 2: Screenshot_from_2024-05-28_16-15-43.png
Screenshot_from_2024-05-28_16-15-43.png
Attachment 3: Screenshot_from_2024-05-28_16-16-18.png
Screenshot_from_2024-05-28_16-16-18.png
Attachment 4: Screenshot_from_2024-05-28_16-20-05.png
Screenshot_from_2024-05-28_16-20-05.png
Attachment 5: Screenshot_from_2024-05-28_16-28-46.png
Screenshot_from_2024-05-28_16-28-46.png
  554   Tue May 28 12:38:31 2024 TD, JM, CBRun 2 - Alpha background run
13.35 FEE64 temperatures OK - attachment 1

      ADC data item rates OK - attachment 2
       overall rate dominated by 2x hot channels aida15

     per FEE64 Stat spectra - attachment 3
      spectra consistent with April alpha background test 
      note high HEC data stats for aida12 asic#2 - can probably be fixed by ASIC control check

     per 1.8.W spectra - 20us FSR - attachments 4-5
      note 100kHz ripple observed for some FEE64s cf. https://elog.ph.ed.ac.uk/CARME/544

Run stopped 15:15 
Attachment 1: Screenshot_from_2024-05-28_13-35-03.png
Screenshot_from_2024-05-28_13-35-03.png
Attachment 2: Screenshot_from_2024-05-28_13-34-30.png
Screenshot_from_2024-05-28_13-34-30.png
Attachment 3: Screenshot_from_2024-05-28_13-33-19.png
Screenshot_from_2024-05-28_13-33-19.png
Attachment 4: Screenshot_from_2024-05-28_13-36-30.png
Screenshot_from_2024-05-28_13-36-30.png
Attachment 5: Screenshot_from_2024-05-28_13-37-16.png
Screenshot_from_2024-05-28_13-37-16.png
  553   Mon May 27 15:25:08 2024 JMStartup

27/05/24

The carme-gsi computer has been moved from the server room into the shifter control room.

Startup of the DAQ looks OK. Settings loaded 2024Apr24

System wide checks OK. Aida16 clock status fails. Aida16 ADC calibration fails. Rest OK.

All FEEs look OK, we again see high rates in aida15 ~ 12k data items/sec, rate in other FEEs is 0.

Testing of the pulser widths has been done. Pulser widths will be uploaded soon.

Background run started. No beam, no target. File G22-201 and R2. Data being written is ~128 kb/s. Run will be left overnight, no night shift.

28/05/24

DAQ continues with alpha background run.

System wide checks OK. Aida16 clock status fails. Aida16 ADC calibration fails. Rest OK. Temps OK

Data can be seen in the spectrum browser on MIDAS. CARME monitoring code was started incorrectly yesterday. It has now be restarted and a small number of events are seen

  552   Fri May 10 21:23:35 2024 TD, CB, JM20Ne7+ test beam

22.00

Carsten Brandau stopping e-target experiment Unattended beam overnight. No operators watching.

10^5 (nominal) 20Ne7+ at 380keV/u beam available

Carsten reports beam is unstable, often disappearing, and in bunched mode.

4x MSL type TTT14 DSSSDs not moving, locked in out position

Starting CARME MIDAS DAQ ASIC settings 2024Apr24-14.14.03

slow comparator 0x64 E

G&G Ortec 448 Pulser OFF

DSSSD bias & leakage currents - attachment 1

FEE64 temperatures OK - attachment 2

All system wide checks OK, WR timestamps OK - attachment 3

ADC data item stats - attachment 4

Rates dominated by two hot channels aida15 per FEE64 Stat spectra - attachment 5

per FEE64 1.8.L spectra - attachments 6-7 no obvious beam energy peak, continuum of energies from below beam energy to threshold

per FEE64 1.8.W spectra - attachments 8-9

 

22.55 Start data file APR24/R5 Merger, TapeServer etc - attachments 10-12

270k data items/s

 

23:29 Activated CARMESort script. See attachments 13-15.

No obvious beam or signal of any kind. Temperatures OK.

 

00:13 DAQ continues OK. System-wide check OK. Temps OK. Still no clear sign of beam.

 

01:03 Daq continues OK. System wide checks OK, Temps OK.

Only low energies observed in the energy spectrum. Looks a bit like the low energy noise observed in our previous runs with no/very few counts above what would be beam energy (~7.6 MeV)

Note the xy plots are not all orientated correctly (strips are correct in the sort code but root axis need to be flipped - I can;t remember how atm). The part of the plots with counts are those closest to the beam. 

02:00 DAQ continues OK. System wide checks OK. Temps OK.

Spectra still looks the same. I have noticed the right hand side rate is higher than the left. EyEx cut spectrum also looks like much more counts outside of the cut.

03:40

DAQ continues. All checks and temps OK.

I can now see a small peak at 5 MeV in the bottom left dssd. This dssd also has the least number of counts in the low energy continuum.

I have logged back into grafana on x2go to get the number ions in the ring. Attachment 17

04:23 (05:23 EU)

DAQ stopped, merger stopped, tape server stopped. Bias OFF. FEEs turned OFF. Screen shots taken of the monitoring code plots - not yet uploaded.

Windows and servers remain up in case we want to do some more tests in the near future.

Attachment 18 shows beam intensity for total run time from Grafana

Attachment 1: Screenshot_from_2024-05-10_22-31-48.png
Screenshot_from_2024-05-10_22-31-48.png
Attachment 2: Screenshot_from_2024-05-10_22-32-52.png
Screenshot_from_2024-05-10_22-32-52.png
Attachment 3: Screenshot_from_2024-05-10_22-33-45.png
Screenshot_from_2024-05-10_22-33-45.png
Attachment 4: Screenshot_from_2024-05-10_22-34-59.png
Screenshot_from_2024-05-10_22-34-59.png
Attachment 5: Screenshot_from_2024-05-10_22-37-12.png
Screenshot_from_2024-05-10_22-37-12.png
Attachment 6: Screenshot_from_2024-05-10_22-38-32.png
Screenshot_from_2024-05-10_22-38-32.png
Attachment 7: Screenshot_from_2024-05-10_22-39-48.png
Screenshot_from_2024-05-10_22-39-48.png
Attachment 8: Screenshot_from_2024-05-10_22-40-38.png
Screenshot_from_2024-05-10_22-40-38.png
Attachment 9: Screenshot_from_2024-05-10_22-41-24.png
Screenshot_from_2024-05-10_22-41-24.png
Attachment 10: Screenshot_from_2024-05-10_22-56-53.png
Screenshot_from_2024-05-10_22-56-53.png
Attachment 11: Screenshot_from_2024-05-10_22-56-20.png
Screenshot_from_2024-05-10_22-56-20.png
Attachment 12: Screenshot_from_2024-05-10_22-56-35.png
Screenshot_from_2024-05-10_22-56-35.png
Attachment 13: Screenshot_from_2024-05-10_23-28-37.png
Screenshot_from_2024-05-10_23-28-37.png
Attachment 14: Screenshot_from_2024-05-10_23-28-45.png
Screenshot_from_2024-05-10_23-28-45.png
Attachment 15: Screenshot_from_2024-05-10_23-28-51.png
Screenshot_from_2024-05-10_23-28-51.png
Attachment 16: Screenshot_from_2024-05-11_00-12-43.png
Screenshot_from_2024-05-11_00-12-43.png
Attachment 17: 20240511_025056.jpg
20240511_025056.jpg
Attachment 18: 20240511_042236.jpg
20240511_042236.jpg
  550   Sun Apr 28 23:12:11 2024 TDMonday 28 April
00.09 DAQ continues data file R3_21

      DSSSD bias & leakage current OK - attachment 1

      FEE64 temps OK - attachment 2

      ADC data item stats OK - attachment 3

      per FEE64 Rate spectra - attachment 4

      Merger, TapeServer etc - attachments 5-7

04.50 DAQ continues data file R3_22

      DSSSD bias & leakage current OK - attachment 8

      FEE64 temps OK - attachment 9

      ADC data item stats OK - attachment 10

      per FEE64 Rate spectra - attachment 11

      Merger, TapeServer etc - attachments 12-14


09.41 DAQ continues data file R3_23

      DSSSD bias & leakage current OK - attachment 15

      FEE64 temps OK - attachment 16

      ADC data item stats OK - attachment 17

      per FEE64 Rate spectra - attachment 18

      Merger, TapeServer etc - attachments 19-21


09.43 DAQ stopped
      DSSSD bias OFF
      FEE64 power OFF
Attachment 1: Screenshot_from_2024-04-29_00-09-18.png
Screenshot_from_2024-04-29_00-09-18.png
Attachment 2: Screenshot_from_2024-04-29_00-09-45.png
Screenshot_from_2024-04-29_00-09-45.png
Attachment 3: Screenshot_from_2024-04-29_00-10-33.png
Screenshot_from_2024-04-29_00-10-33.png
Attachment 4: Screenshot_from_2024-04-29_00-10-52.png
Screenshot_from_2024-04-29_00-10-52.png
Attachment 5: Screenshot_from_2024-04-29_00-11-04.png
Screenshot_from_2024-04-29_00-11-04.png
Attachment 6: Screenshot_from_2024-04-29_00-11-29.png
Screenshot_from_2024-04-29_00-11-29.png
Attachment 7: Screenshot_from_2024-04-29_00-11-47.png
Screenshot_from_2024-04-29_00-11-47.png
Attachment 8: Screenshot_from_2024-04-29_04-50-25.png
Screenshot_from_2024-04-29_04-50-25.png
Attachment 9: Screenshot_from_2024-04-29_04-50-52.png
Screenshot_from_2024-04-29_04-50-52.png
Attachment 10: Screenshot_from_2024-04-29_04-51-38.png
Screenshot_from_2024-04-29_04-51-38.png
Attachment 11: Screenshot_from_2024-04-29_04-52-01.png
Screenshot_from_2024-04-29_04-52-01.png
Attachment 12: Screenshot_from_2024-04-29_04-52-43.png
Screenshot_from_2024-04-29_04-52-43.png
Attachment 13: Screenshot_from_2024-04-29_04-53-05.png
Screenshot_from_2024-04-29_04-53-05.png
Attachment 14: Screenshot_from_2024-04-29_04-53-22.png
Screenshot_from_2024-04-29_04-53-22.png
Attachment 15: Screenshot_from_2024-04-29_09-41-28.png
Screenshot_from_2024-04-29_09-41-28.png
Attachment 16: Screenshot_from_2024-04-29_09-42-01.png
Screenshot_from_2024-04-29_09-42-01.png
Attachment 17: Screenshot_from_2024-04-29_09-42-43.png
Screenshot_from_2024-04-29_09-42-43.png
Attachment 18: Screenshot_from_2024-04-29_09-42-58.png
Screenshot_from_2024-04-29_09-42-58.png
Attachment 19: Screenshot_from_2024-04-29_09-43-08.png
Screenshot_from_2024-04-29_09-43-08.png
Attachment 20: Screenshot_from_2024-04-29_09-43-28.png
Screenshot_from_2024-04-29_09-43-28.png
Attachment 21: Screenshot_from_2024-04-29_09-43-41.png
Screenshot_from_2024-04-29_09-43-41.png
  549   Sat Apr 27 23:16:02 2024 TDSunday 28 April
00.12 DAQ continues data file R3_15

      DSSSD bias & leakage current OK - attachment 1

      FEE64 temps OK - attachment 2

      ADC data item stats OK - attachment 3

      per FEE64 Rate spectra - attachment 4

      Merger, TapeServer etc - attachments 5-7


05.32 DAQ continues data file R3_16

      DSSSD bias & leakage current OK - attachment 8

      FEE64 temps OK - attachment 9

      ADC data item stats OK - attachment 10

      per FEE64 Rate spectra - attachment 11

      Merger, TapeServer etc - attachments 12-14

11.27 DAQ continues data file R3_17

      DSSSD bias & leakage current OK - attachment 15

      FEE64 temps OK - attachment 16

      ADC data item stats OK - attachment 17

      per FEE64 Rate spectra - attachment 21

      Merger, TapeServer etc - attachments 18-20

11.33 aida16 asic #2 2x Check ASIC control
      aida16 asic #4 2x Check ASIC control


18.53 DAQ continues data file R3_19

      DSSSD bias & leakage current OK - attachment 22

      FEE64 temps OK - attachment 23

      ADC data item stats OK - attachment 24

      per FEE64 Rate spectra - attachment 25

      Merger, TapeServer etc - attachments 26-28

      per FEE64 1.8.H - attachments 29-30

      per FEE64 1.8.L - attachments 31-32

      per FEE64 1.8.W - 20us FSR - attachments 33-36

19.10 CRYRING ambient temperature per https://web-docs.gsi.de/~lestinsk/tempplot.php - attachment 37
Attachment 1: Screenshot_from_2024-04-28_00-11-59.png
Screenshot_from_2024-04-28_00-11-59.png
Attachment 2: Screenshot_from_2024-04-28_00-12-25.png
Screenshot_from_2024-04-28_00-12-25.png
Attachment 3: Screenshot_from_2024-04-28_00-13-20.png
Screenshot_from_2024-04-28_00-13-20.png
Attachment 4: Screenshot_from_2024-04-28_00-14-35.png
Screenshot_from_2024-04-28_00-14-35.png
Attachment 5: Screenshot_from_2024-04-28_00-15-03.png
Screenshot_from_2024-04-28_00-15-03.png
Attachment 6: Screenshot_from_2024-04-28_00-15-40.png
Screenshot_from_2024-04-28_00-15-40.png
Attachment 7: Screenshot_from_2024-04-28_00-15-20.png
Screenshot_from_2024-04-28_00-15-20.png
Attachment 8: Screenshot_from_2024-04-28_05-32-48.png
Screenshot_from_2024-04-28_05-32-48.png
Attachment 9: Screenshot_from_2024-04-28_05-33-11.png
Screenshot_from_2024-04-28_05-33-11.png
Attachment 10: Screenshot_from_2024-04-28_05-34-03.png
Screenshot_from_2024-04-28_05-34-03.png
Attachment 11: Screenshot_from_2024-04-28_05-34-36.png
Screenshot_from_2024-04-28_05-34-36.png
Attachment 12: Screenshot_from_2024-04-28_05-34-50.png
Screenshot_from_2024-04-28_05-34-50.png
Attachment 13: Screenshot_from_2024-04-28_05-35-08.png
Screenshot_from_2024-04-28_05-35-08.png
Attachment 14: Screenshot_from_2024-04-28_05-35-22.png
Screenshot_from_2024-04-28_05-35-22.png
Attachment 15: Screenshot_from_2024-04-28_11-27-15.png
Screenshot_from_2024-04-28_11-27-15.png
Attachment 16: Screenshot_from_2024-04-28_11-27-35.png
Screenshot_from_2024-04-28_11-27-35.png
Attachment 17: Screenshot_from_2024-04-28_11-27-52.png
Screenshot_from_2024-04-28_11-27-52.png
Attachment 18: Screenshot_from_2024-04-28_11-28-48.png
Screenshot_from_2024-04-28_11-28-48.png
Attachment 19: Screenshot_from_2024-04-28_11-29-02.png
Screenshot_from_2024-04-28_11-29-02.png
Attachment 20: Screenshot_from_2024-04-28_11-29-13.png
Screenshot_from_2024-04-28_11-29-13.png
Attachment 21: Screenshot_from_2024-04-28_11-31-05.png
Screenshot_from_2024-04-28_11-31-05.png
Attachment 22: Screenshot_from_2024-04-28_18-53-09.png
Screenshot_from_2024-04-28_18-53-09.png
Attachment 23: Screenshot_from_2024-04-28_18-53-32.png
Screenshot_from_2024-04-28_18-53-32.png
Attachment 24: Screenshot_from_2024-04-28_18-53-54.png
Screenshot_from_2024-04-28_18-53-54.png
Attachment 25: Screenshot_from_2024-04-28_18-54-40.png
Screenshot_from_2024-04-28_18-54-40.png
Attachment 26: Screenshot_from_2024-04-28_18-54-50.png
Screenshot_from_2024-04-28_18-54-50.png
Attachment 27: Screenshot_from_2024-04-28_18-55-06.png
Screenshot_from_2024-04-28_18-55-06.png
Attachment 28: Screenshot_from_2024-04-28_18-55-17.png
Screenshot_from_2024-04-28_18-55-17.png
Attachment 29: Screenshot_from_2024-04-28_19-00-46.png
Screenshot_from_2024-04-28_19-00-46.png
Attachment 30: Screenshot_from_2024-04-28_19-01-27.png
Screenshot_from_2024-04-28_19-01-27.png
Attachment 31: Screenshot_from_2024-04-28_18-59-12.png
Screenshot_from_2024-04-28_18-59-12.png
Attachment 32: Screenshot_from_2024-04-28_19-00-08.png
Screenshot_from_2024-04-28_19-00-08.png
Attachment 33: Screenshot_from_2024-04-28_19-01-57.png
Screenshot_from_2024-04-28_19-01-57.png
Attachment 34: Screenshot_from_2024-04-28_19-02-21.png
Screenshot_from_2024-04-28_19-02-21.png
Attachment 35: Screenshot_from_2024-04-28_19-02-53.png
Screenshot_from_2024-04-28_19-02-53.png
Attachment 36: Screenshot_from_2024-04-28_19-03-18.png
Screenshot_from_2024-04-28_19-03-18.png
Attachment 37: CRY_temps.PNG
CRY_temps.PNG
  548   Fri Apr 26 23:41:16 2024 TDSaturday 27 April
00.34 DAQ continues data file R3_9

      DSSSD bias & leakage current OK - attachment 1

      FEE64 temps OK - attachment 2

      ADC data item stats OK - attachment 3

      per FEE64 Rate spectra - attachment 4

      Merger, TapeServer etc - attachments 5-7

06.41 DAQ continues data file R3_10

      DSSSD bias & leakage current OK - attachment 8

      FEE64 temps OK - attachment 9

      ADC data item stats OK - attachment 10

      per FEE64 Rate spectra - attachment 11

      Merger, TapeServer etc - attachments 12-14


10.56 DAQ continues data file R3_11

      DSSSD bias & leakage current OK - attachment 15

      FEE64 temps OK - attachment 16

      ADC data item stats OK - attachment 17

      per FEE64 Rate spectra - attachment 18

      Merger, TapeServer etc - attachments 19-21

14.33 DAQ continues data file R3_12

      DSSSD bias & leakage current OK - attachment 22

      FEE64 temps OK - attachment 23

      ADC data item stats OK - attachment 24

      per FEE64 Rate spectra - attachment 25

      Merger, TapeServer etc - attachments 26-28

18.31 DAQ continues data file R3_13

      DSSSD bias & leakage current OK - attachment 29

      FEE64 temps OK - attachment 30

      ADC data item stats OK - attachment 31

      per FEE64 Rate spectra - attachment 32

      per FEE64 1.8.L spectra - attachments 33-34

      Merger, TapeServer etc - attachments 35-37
Attachment 1: Screenshot_from_2024-04-27_00-34-52.png
Screenshot_from_2024-04-27_00-34-52.png
Attachment 2: Screenshot_from_2024-04-27_00-35-25.png
Screenshot_from_2024-04-27_00-35-25.png
Attachment 3: Screenshot_from_2024-04-27_00-36-07.png
Screenshot_from_2024-04-27_00-36-07.png
Attachment 4: Screenshot_from_2024-04-27_00-36-34.png
Screenshot_from_2024-04-27_00-36-34.png
Attachment 5: Screenshot_from_2024-04-27_00-36-49.png
Screenshot_from_2024-04-27_00-36-49.png
Attachment 6: Screenshot_from_2024-04-27_00-37-14.png
Screenshot_from_2024-04-27_00-37-14.png
Attachment 7: Screenshot_from_2024-04-27_00-37-37.png
Screenshot_from_2024-04-27_00-37-37.png
Attachment 8: Screenshot_from_2024-04-27_06-41-34.png
Screenshot_from_2024-04-27_06-41-34.png
Attachment 9: Screenshot_from_2024-04-27_06-42-06.png
Screenshot_from_2024-04-27_06-42-06.png
Attachment 10: Screenshot_from_2024-04-27_06-42-48.png
Screenshot_from_2024-04-27_06-42-48.png
Attachment 11: Screenshot_from_2024-04-27_06-43-08.png
Screenshot_from_2024-04-27_06-43-08.png
Attachment 12: Screenshot_from_2024-04-27_06-43-19.png
Screenshot_from_2024-04-27_06-43-19.png
Attachment 13: Screenshot_from_2024-04-27_06-43-36.png
Screenshot_from_2024-04-27_06-43-36.png
Attachment 14: Screenshot_from_2024-04-27_06-43-50.png
Screenshot_from_2024-04-27_06-43-50.png
Attachment 15: Screenshot_from_2024-04-27_10-55-54.png
Screenshot_from_2024-04-27_10-55-54.png
Attachment 16: Screenshot_from_2024-04-27_10-56-21.png
Screenshot_from_2024-04-27_10-56-21.png
Attachment 17: Screenshot_from_2024-04-27_10-56-38.png
Screenshot_from_2024-04-27_10-56-38.png
Attachment 18: Screenshot_from_2024-04-27_10-57-26.png
Screenshot_from_2024-04-27_10-57-26.png
Attachment 19: Screenshot_from_2024-04-27_10-57-45.png
Screenshot_from_2024-04-27_10-57-45.png
Attachment 20: Screenshot_from_2024-04-27_10-58-04.png
Screenshot_from_2024-04-27_10-58-04.png
Attachment 21: Screenshot_from_2024-04-27_10-58-16.png
Screenshot_from_2024-04-27_10-58-16.png
Attachment 22: Screenshot_from_2024-04-27_14-33-33.png
Screenshot_from_2024-04-27_14-33-33.png
Attachment 23: Screenshot_from_2024-04-27_14-34-04.png
Screenshot_from_2024-04-27_14-34-04.png
Attachment 24: Screenshot_from_2024-04-27_14-34-52.png
Screenshot_from_2024-04-27_14-34-52.png
Attachment 25: Screenshot_from_2024-04-27_14-35-08.png
Screenshot_from_2024-04-27_14-35-08.png
Attachment 26: Screenshot_from_2024-04-27_14-35-19.png
Screenshot_from_2024-04-27_14-35-19.png
Attachment 27: Screenshot_from_2024-04-27_14-35-50.png
Screenshot_from_2024-04-27_14-35-50.png
Attachment 28: Screenshot_from_2024-04-27_14-35-34.png
Screenshot_from_2024-04-27_14-35-34.png
Attachment 29: Screenshot_from_2024-04-27_18-31-01.png
Screenshot_from_2024-04-27_18-31-01.png
Attachment 30: Screenshot_from_2024-04-27_18-31-22.png
Screenshot_from_2024-04-27_18-31-22.png
Attachment 31: Screenshot_from_2024-04-27_18-31-40.png
Screenshot_from_2024-04-27_18-31-40.png
Attachment 32: Screenshot_from_2024-04-27_18-32-27.png
Screenshot_from_2024-04-27_18-32-27.png
Attachment 33: Screenshot_from_2024-04-27_18-33-04.png
Screenshot_from_2024-04-27_18-33-04.png
Attachment 34: Screenshot_from_2024-04-27_18-33-47.png
Screenshot_from_2024-04-27_18-33-47.png
Attachment 35: Screenshot_from_2024-04-27_18-34-04.png
Screenshot_from_2024-04-27_18-34-04.png
Attachment 36: Screenshot_from_2024-04-27_18-34-22.png
Screenshot_from_2024-04-27_18-34-22.png
Attachment 37: Screenshot_from_2024-04-27_18-34-44.png
Screenshot_from_2024-04-27_18-34-44.png
  547   Fri Apr 26 07:05:01 2024 TDFriday 26 April
08.02 DAQ continues data file R1_14

      DSSSD bias & leakage current OK - attachment 1

      FEE64 temps OK - attachment 2

      ADC data item stats OK - attachment 3

      per FEE64 Rate spectra - attachment 4 

      Merger, TapeServer etc - attachments 5-7

10.40 Close file R1 and open file R2

      Switch on pulser (relay line 50Hz)

      per FEE64 Rate spectra - attachment 8

      Check ASIC load x2 for aida09, aida10, aida12, aida15  aida16

      per FEE64 Rate spectra - attachment 12
       status quo ante

      Pulser walkthrough 

      Pulser settings per https://elog.ph.ed.ac.uk/CARME/544
      Amplituides 70,000-10,000 @ 10,000 step

      per FEE64 1.8.L spectra - attachments 9-10
 
      per FEE64 Rate spectra - attachment 11

11.05 Pulser OFF
      Close file R2, open file R3

      zero all histograms  stats

      background alpha run continues


20.33 DAQ continues data file R3_8

      DSSSD bias & leakage current OK - attachment 13

      FEE64 temps OK - attachment 14

      ADC data item stats OK - attachment 15

      per FEE64 Rate spectra - attachment 16

      Merger, TapeServer etc - attachments 1-19


      CAEN data log stopped 00:21 Friday 26 April 

pi@nnrpi1:~/Programs/caenlogger $ caenlogger --config caenlogger_config_gsi.toml
2024-04-25 15:25:16.194 | INFO     | caenlogger.__main__:main:117 - Config file provided.
Traceback (most recent call last):
  File "/home/pi/.local/bin/caenlogger", line 10, in <module>
    sys.exit(main())
  File "/home/pi/.local/lib/python3.7/site-packages/caenlogger/__main__.py", line 135, in main
    curses.wrapper(process, config_dic)
  File "/usr/lib/python3.7/curses/__init__.py", line 94, in wrapper
    return func(stdscr, *args, **kwds)
  File "/home/pi/.local/lib/python3.7/site-packages/caenlogger/__main__.py", line 57, in process
    Vmon = [ f for f in HV[5].strip().split("  ") if f ]
IndexError: list index out of range

     Old log gzip'd and moved to directory ~/logs

[npg@carme-gsi ~]$ mv caen_current.dat.gz logs/0031260424-caen_current.dat.gz

20.45 Restarted logger program
Attachment 1: Screenshot_from_2024-04-26_08-02-37.png
Screenshot_from_2024-04-26_08-02-37.png
Attachment 2: Screenshot_from_2024-04-26_08-02-57.png
Screenshot_from_2024-04-26_08-02-57.png
Attachment 3: Screenshot_from_2024-04-26_08-03-17.png
Screenshot_from_2024-04-26_08-03-17.png
Attachment 4: Screenshot_from_2024-04-26_08-03-56.png
Screenshot_from_2024-04-26_08-03-56.png
Attachment 5: Screenshot_from_2024-04-26_08-04-06.png
Screenshot_from_2024-04-26_08-04-06.png
Attachment 6: Screenshot_from_2024-04-26_08-04-25.png
Screenshot_from_2024-04-26_08-04-25.png
Attachment 7: Screenshot_from_2024-04-26_08-04-41.png
Screenshot_from_2024-04-26_08-04-41.png
Attachment 8: Screenshot_from_2024-04-26_10-38-00.png
Screenshot_from_2024-04-26_10-38-00.png
Attachment 9: Screenshot_from_2024-04-26_11-01-53.png
Screenshot_from_2024-04-26_11-01-53.png
Attachment 10: Screenshot_from_2024-04-26_11-02-51.png
Screenshot_from_2024-04-26_11-02-51.png
Attachment 11: Screenshot_from_2024-04-26_11-03-23.png
Screenshot_from_2024-04-26_11-03-23.png
Attachment 12: Screenshot_from_2024-04-26_10-43-19.png
Screenshot_from_2024-04-26_10-43-19.png
Attachment 13: Screenshot_from_2024-04-26_20-33-24.png
Screenshot_from_2024-04-26_20-33-24.png
Attachment 14: Screenshot_from_2024-04-26_20-33-49.png
Screenshot_from_2024-04-26_20-33-49.png
Attachment 15: Screenshot_from_2024-04-26_20-34-30.png
Screenshot_from_2024-04-26_20-34-30.png
Attachment 16: Screenshot_from_2024-04-26_20-36-04.png
Screenshot_from_2024-04-26_20-36-04.png
Attachment 17: Screenshot_from_2024-04-26_20-36-17.png
Screenshot_from_2024-04-26_20-36-17.png
Attachment 18: Screenshot_from_2024-04-26_20-36-37.png
Screenshot_from_2024-04-26_20-36-37.png
Attachment 19: Screenshot_from_2024-04-26_20-36-49.png
Screenshot_from_2024-04-26_20-36-49.png
  546   Thu Apr 25 02:31:25 2024 TDThursday 25 April
03.27 DAQ continues - file APR24/R1_7

      DSSSD bias & leakage current OK - attachment 1
      ambient temperature 16.1 deg C

      FEE64 temperatures OK - attachment 2

      ADC data item stats OK - attachment 3

      per n+n FEE64 1.8.L spectra - attachment 4

      per FEE64 Stat spectra - attachment 5

      Merger, TapeServer etc - attachments 6-8


10.19 DAQ continues - file APR24/R1_8

      DSSSD bias & leakage current OK - attachment 9
      ambient temperature 16.2 deg C

      FEE64 temperatures OK - attachment 10

      ADC data item stats OK - attachment 11

      per FEE64 Stat spectra - attachment 12

      Merger, TapeServer etc - attachments 13-15

      NetVar merge variables - attachment 16


14.44 DAQ continues - file APR24/R1_10

      DSSSD bias & leakage current OK - attachment 17
      ambient temperature 16.5 deg C

      FEE64 temperatures OK - attachment 18

      ADC data item stats OK - attachment 19

      per FEE64 Stat spectra - attachment 20
       some HEC data has appeared

      Merger, TapeServer etc - attachments 21-23

19.12 DAQ continues - file APR24/R1_11

      DSSSD bias & leakage current OK - attachment 24
      ambient temperature 16.5 deg C

      FEE64 temperatures OK - attachment 25

      ADC data item stats OK - attachment 26

      per FEE64 Stat spectra - attachment 27

      per FEE64 1.8.L spectra - attachments 28-29

      Merger, TapeServer etc - attachments 30-32

23.14 DAQ continues - file APR24/R1_12

      DSSSD bias & leakage current OK - attachment 33
      ambient temperature 16.1 deg C

      FEE64 temperatures OK - attachment 34

      ADC data item stats OK - attachment 35

      per FEE64 Stat spectra - attachment 36

      Merger, TapeServer etc - attachments 37-39
Attachment 1: Screenshot_from_2024-04-25_03-26-58.png
Screenshot_from_2024-04-25_03-26-58.png
Attachment 2: Screenshot_from_2024-04-25_03-27-48.png
Screenshot_from_2024-04-25_03-27-48.png
Attachment 3: Screenshot_from_2024-04-25_03-28-19.png
Screenshot_from_2024-04-25_03-28-19.png
Attachment 4: Screenshot_from_2024-04-25_03-28-41.png
Screenshot_from_2024-04-25_03-28-41.png
Attachment 5: Screenshot_from_2024-04-25_03-32-43.png
Screenshot_from_2024-04-25_03-32-43.png
Attachment 6: Screenshot_from_2024-04-25_03-30-13.png
Screenshot_from_2024-04-25_03-30-13.png
Attachment 7: Screenshot_from_2024-04-25_03-30-37.png
Screenshot_from_2024-04-25_03-30-37.png
Attachment 8: Screenshot_from_2024-04-25_03-30-55.png
Screenshot_from_2024-04-25_03-30-55.png
Attachment 9: Screenshot_from_2024-04-25_10-19-09.png
Screenshot_from_2024-04-25_10-19-09.png
Attachment 10: Screenshot_from_2024-04-25_10-19-31.png
Screenshot_from_2024-04-25_10-19-31.png
Attachment 11: Screenshot_from_2024-04-25_10-19-54.png
Screenshot_from_2024-04-25_10-19-54.png
Attachment 12: Screenshot_from_2024-04-25_10-20-14.png
Screenshot_from_2024-04-25_10-20-14.png
Attachment 13: Screenshot_from_2024-04-25_10-20-26.png
Screenshot_from_2024-04-25_10-20-26.png
Attachment 14: Screenshot_from_2024-04-25_10-20-53.png
Screenshot_from_2024-04-25_10-20-53.png
Attachment 15: Screenshot_from_2024-04-25_10-21-22.png
Screenshot_from_2024-04-25_10-21-22.png
Attachment 16: Screenshot_from_2024-04-25_10-21-45.png
Screenshot_from_2024-04-25_10-21-45.png
Attachment 17: Screenshot_from_2024-04-25_14-44-05.png
Screenshot_from_2024-04-25_14-44-05.png
Attachment 18: Screenshot_from_2024-04-25_14-44-29.png
Screenshot_from_2024-04-25_14-44-29.png
Attachment 19: Screenshot_from_2024-04-25_14-44-49.png
Screenshot_from_2024-04-25_14-44-49.png
Attachment 20: Screenshot_from_2024-04-25_14-45-51.png
Screenshot_from_2024-04-25_14-45-51.png
Attachment 21: Screenshot_from_2024-04-25_14-46-35.png
Screenshot_from_2024-04-25_14-46-35.png
Attachment 22: Screenshot_from_2024-04-25_14-46-53.png
Screenshot_from_2024-04-25_14-46-53.png
Attachment 23: Screenshot_from_2024-04-25_14-47-05.png
Screenshot_from_2024-04-25_14-47-05.png
Attachment 24: Screenshot_from_2024-04-25_19-12-30.png
Screenshot_from_2024-04-25_19-12-30.png
Attachment 25: Screenshot_from_2024-04-25_19-13-01.png
Screenshot_from_2024-04-25_19-13-01.png
Attachment 26: Screenshot_from_2024-04-25_19-13-39.png
Screenshot_from_2024-04-25_19-13-39.png
Attachment 27: Screenshot_from_2024-04-25_19-14-02.png
Screenshot_from_2024-04-25_19-14-02.png
Attachment 28: Screenshot_from_2024-04-25_19-14-40.png
Screenshot_from_2024-04-25_19-14-40.png
Attachment 29: Screenshot_from_2024-04-25_19-15-24.png
Screenshot_from_2024-04-25_19-15-24.png
Attachment 30: Screenshot_from_2024-04-25_19-15-58.png
Screenshot_from_2024-04-25_19-15-58.png
Attachment 31: Screenshot_from_2024-04-25_19-16-12.png
Screenshot_from_2024-04-25_19-16-12.png
Attachment 32: Screenshot_from_2024-04-25_19-16-30.png
Screenshot_from_2024-04-25_19-16-30.png
Attachment 33: Screenshot_from_2024-04-25_23-14-01.png
Screenshot_from_2024-04-25_23-14-01.png
Attachment 34: Screenshot_from_2024-04-25_23-14-23.png
Screenshot_from_2024-04-25_23-14-23.png
Attachment 35: Screenshot_from_2024-04-25_23-14-41.png
Screenshot_from_2024-04-25_23-14-41.png
Attachment 36: Screenshot_from_2024-04-25_23-16-04.png
Screenshot_from_2024-04-25_23-16-04.png
Attachment 37: Screenshot_from_2024-04-25_23-16-12.png
Screenshot_from_2024-04-25_23-16-12.png
Attachment 38: Screenshot_from_2024-04-25_23-16-25.png
Screenshot_from_2024-04-25_23-16-25.png
Attachment 39: Screenshot_from_2024-04-25_23-16-42.png
Screenshot_from_2024-04-25_23-16-42.png
  545   Wed Apr 24 13:18:09 2024 TDExamples of poor CARME cable management/routing
Have observed some instances of poor CARME cable management/routing
Attachment 1 - SAES D-2000 ion pump HV cable (netal braid) transiting FEE64 cabling
Attachment 2 - CARME motor controller cable (grey plastic) transiting FEE64 cabling

General check and tidying up required.
Attachment 1: 20240424_125858.jpg
20240424_125858.jpg
Attachment 2: 20240424_125929.jpg
20240424_125929.jpg
  544   Wed Apr 24 12:48:35 2024 TDWednesday 24 April
      Replace AIDA ASIC mezzanine of aida01

13.43 DSSSD bias & leakage current OK - attachment 1
      ambient temperature 16.5 deg C

      FEE64 temperatures OK - attachment 2

      All system wide checks OK 

      WR timestamps OK - attachment 3

      ADC data item stats OK - attachment 4

      per FEE64 Rate spectra - attachment 5
       c. 10 strips (of 1024) not producing data
       most missing strips are from aida05 - misaligned or damaged adaptor PCB?

      per p+n FEE64 1.8.L spectra - attachment 6
       aida01 pulser peak width 75 ch FWHM
       common x/y scale => peak height proxy for peak width

      per n+n FEE64 1.8.L spectra - attachment 7
       aida01 pulser peak width 70 ch FWHM

      per FEE64 1.8.W spectra - 20us FSR - attachments 8-11

Summary

All FEE64s appear to be working (except aida05 which has c. 8 missing channels) with pulser - which 
validates the FEE64, ASICs and adaptor PCB.

Electronic noise c. 50keV FWHM for aida01 & aida03 - other FEE64s better/worse.

Preamplifier noise appears to be similar for all FEE64s which may suggest an issue with the test signal chain

Note
1) e-target is starting up
2) have identified some instances of poor CARME cable management/routing see https://elog.ph.ed.ac.uk/CARME/545



EG&G Ortec 448 Pulser
Amplitude 50,000
Normalise 5.00
x5 attenuator IN
t_r 100ns
tau_d 50us
Polarity +
Frequency relay line (50Hz)

Cooknell SA1 Summing Amplifier
inverted output to test - daisy chain

Current ASIC settings saved to 2024Apr24-14.14.03


14.48 Pulser OFF
 
      All histograms & stats zero'd

      LEC slow comparator 0x64 -> 0xff
      HEC fast comparator 0x2 -> 0x5

      background alpha run to validate signal path from DSSSD to FEE64 adaptor PCB

22.18 Start Merger & TapeServer - attachments 12-14

      DSSSD bias & leakage current OK - attachment 21
      ambient temperature 16.0 deg C


      FEE64 temperatures OK - attachment 15

      WR timestamps OK - attachment 16

      ADC data item stats OK - attachment 17

      per FEE64 Rate spectra - attachment 18

      per p+n FEE64 1.8.L spectra - attachments 19-20
Attachment 1: Screenshot_from_2024-04-24_13-43-29.png
Screenshot_from_2024-04-24_13-43-29.png
Attachment 2: Screenshot_from_2024-04-24_13-44-02.png
Screenshot_from_2024-04-24_13-44-02.png
Attachment 3: Screenshot_from_2024-04-24_13-45-01.png
Screenshot_from_2024-04-24_13-45-01.png
Attachment 4: Screenshot_from_2024-04-24_13-45-28.png
Screenshot_from_2024-04-24_13-45-28.png
Attachment 5: Screenshot_from_2024-04-24_13-45-46.png
Screenshot_from_2024-04-24_13-45-46.png
Attachment 6: Screenshot_from_2024-04-24_13-47-59.png
Screenshot_from_2024-04-24_13-47-59.png
Attachment 7: Screenshot_from_2024-04-24_13-52-50.png
Screenshot_from_2024-04-24_13-52-50.png
Attachment 8: Screenshot_from_2024-04-24_13-55-02.png
Screenshot_from_2024-04-24_13-55-02.png
Attachment 9: Screenshot_from_2024-04-24_13-55-32.png
Screenshot_from_2024-04-24_13-55-32.png
Attachment 10: Screenshot_from_2024-04-24_13-56-48.png
Screenshot_from_2024-04-24_13-56-48.png
Attachment 11: Screenshot_from_2024-04-24_13-55-59.png
Screenshot_from_2024-04-24_13-55-59.png
Attachment 12: Screenshot_from_2024-04-24_22-19-55.png
Screenshot_from_2024-04-24_22-19-55.png
Attachment 13: Screenshot_from_2024-04-24_22-20-16.png
Screenshot_from_2024-04-24_22-20-16.png
Attachment 14: Screenshot_from_2024-04-24_22-20-30.png
Screenshot_from_2024-04-24_22-20-30.png
Attachment 15: Screenshot_from_2024-04-24_22-21-58.png
Screenshot_from_2024-04-24_22-21-58.png
Attachment 16: Screenshot_from_2024-04-24_22-22-22.png
Screenshot_from_2024-04-24_22-22-22.png
Attachment 17: Screenshot_from_2024-04-24_22-22-57.png
Screenshot_from_2024-04-24_22-22-57.png
Attachment 18: Screenshot_from_2024-04-24_22-23-27.png
Screenshot_from_2024-04-24_22-23-27.png
Attachment 19: Screenshot_from_2024-04-24_22-24-37.png
Screenshot_from_2024-04-24_22-24-37.png
Attachment 20: Screenshot_from_2024-04-24_22-25-30.png
Screenshot_from_2024-04-24_22-25-30.png
Attachment 21: Screenshot_from_2024-04-24_22-27-58.png
Screenshot_from_2024-04-24_22-27-58.png
  543   Tue Apr 23 17:44:41 2024 TDTuesday 23 April
18.43 ASIC settings 2024Feb08-15.59.09

aida06 pulser peak width 66 ch FWHM
Attachment 1: Screenshot_from_2024-04-23_18-53-39.png
Screenshot_from_2024-04-23_18-53-39.png
Attachment 2: Screenshot_from_2024-04-23_18-56-31.png
Screenshot_from_2024-04-23_18-56-31.png
Attachment 3: Screenshot_from_2024-04-23_18-58-49.png
Screenshot_from_2024-04-23_18-58-49.png
Attachment 4: Screenshot_from_2024-04-23_18-59-18.png
Screenshot_from_2024-04-23_18-59-18.png
Attachment 5: Screenshot_from_2024-04-23_19-02-02.png
Screenshot_from_2024-04-23_19-02-02.png
Attachment 6: Screenshot_from_2024-04-23_19-05-12.png
Screenshot_from_2024-04-23_19-05-12.png
Attachment 7: Screenshot_from_2024-04-23_19-05-42.png
Screenshot_from_2024-04-23_19-05-42.png
Attachment 8: Screenshot_from_2024-04-23_19-06-15.png
Screenshot_from_2024-04-23_19-06-15.png
Attachment 9: Screenshot_from_2024-04-23_19-10-40.png
Screenshot_from_2024-04-23_19-10-40.png
Attachment 10: Screenshot_from_2024-04-23_19-14-23.png
Screenshot_from_2024-04-23_19-14-23.png
Attachment 11: Screenshot_from_2024-04-23_19-16-12.png
Screenshot_from_2024-04-23_19-16-12.png
Attachment 12: Screenshot_from_2024-04-23_19-16-46.png
Screenshot_from_2024-04-23_19-16-46.png
ELOG V3.1.4-unknown